OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

\[ \|\mathbf u \cdot \mathbf v\| = \|\mathbf u \| \|\mathbf v\|\cos(\theta) \]

OpenStudy (anonymous):

\[ \theta = \cos^{-1}\left(\frac{\|\mathbf u \cdot \mathbf v\|}{\| \mathbf u\|\|\mathbf v\|}\right) \]

OpenStudy (anonymous):

This is one way of doing it.

OpenStudy (anonymous):

\[ \tan\theta _{\mathbf u} =\frac { \mathbf u_y }{\mathbf u_x} \]

OpenStudy (anonymous):

\[ \theta = |\theta_{\mathbf u} - \theta_{\mathbf v}| \]

OpenStudy (anonymous):

I got -2.

OpenStudy (anonymous):

It's off because these are my choices: -9.1° 1.8° 0.9° 11.8°